# Algorithmic Primitives and Compositional Geometry of Reasoning in Language Models

### v1

@@ Abstract
How do latent and inference time computations enable large language models (LLMs) to solve multi-step reasoning? We introduce a framework for tracing and steering algorithmic primitives that underlie model reasoning. Our approach links reasoning traces to internal activation patterns and evaluates algorithmic primitives by injecting them into residual streams and measuring their effect on reasoning steps and task performance. We consider four benchmarks: Traveling Salesperson Problem (TSP), 3SAT, AIME, and graph navigation. We operationalize primitives by clustering neural activations and labeling their matched reasoning traces. We then apply function vector methods to derive primitive vectors as reusable compositional building blocks of reasoning. Primitive vectors can be combined through addition, subtraction, and scalar operations, revealing a geometric logic in activation space. Cross-task and cross-model evaluations (Phi-4, Phi-4-Reasoning, Llama-3-8B) show both shared and task-specific primitives. Notably, comparing Phi-4 with its reasoning-finetuned variant highlights compositional generalization after finetuning: Phi-4-Reasoning exhibits more systematic use of verification and path-generation primitives. Injecting the associated primitive vectors in Phi-4-Base induces behavioral hallmarks associated with Phi-4-Reasoning. Together, these findings demonstrate that reasoning in LLMs may be supported by a compositional geometry of algorithmic primitives, that primitives transfer cross-task and cross-model, and that reasoning finetuning strengthens algorithmic generalization across domains.
@@ 1 Introduction
This work aims to understand how fundamental algorithmic primitives are generalized and composed to enable complex reasoning capabilities in LLMs. The work addresses three fundamental questions: (1) What are the basic algorithmic primitives that language models use in specific tasks, and across reasoning domains? (2) Do these primitives compose geometrically in neural activation space? (3) How do reasoning-enhanced models differ from base models in their primitive usage and composition? We introduce a framework for a multi-level algorithmic understanding. We first extract algorithmic primitives by clustering internal representations and interpreting the corresponding reasoning traces. We then apply function vector methods to extract primitive vectors from the models’ internal representations. We induce and steer algorithmic primitives by injecting primitive vectors across the layers. To identify both task-specific and universal algorithmic primitives, we apply the approach across domains: Traveling Salesperson Problem, 3-SAT, AIME, and graph navigation. Finally, to investigate compositional generalization as a result of finetuning, we compare algorithmic performance of base models and reasoning-finetuned counterpart, like Phi-4 and Phi4-reasoning (Abdin et al., 2025). This multi-level framework allows us to understand the geometry and compositionality of algorithmic primitives in LLM reasoning.
@@ 2 Related Work
Interpretability for Transformers. With increasing scale and complexity of LLMs, the challenge of understanding their predictions has become an important research direction in explainable AI and interpretability research. This has specifically targeted the analysis of internal model structure and feature relationships (Geiger et al., 2022; Eberle et al., 2022; Schnake et al., 2022), as well as representations and manifolds (Kornblith et al., 2019) including concepts (Chormai et al., 2024). Mechanistic interpretability (Sharkey et al., 2025) has focused on the extraction of circuits (Olah et al., 2020; Wang et al., 2023), feature descriptions (Hernandez et al., 2022), and causally effective representations such as function vectors (Todd et al., 2024). Combined analysis of individual neurons (Gurnee & Tegmark, 2024; Templeton et al., 2024), attention scores (Voita et al., 2019; Clark et al., 2019), and task-specific attention heads (Vig & Belinkov, 2019; McDougall et al., 2024) have further deepened our understanding of internal model processing. In parallel, free-text and chain-of-thought explanations (Turpin et al., 2023; Camburu et al., 2018; Huang et al., 2023; Madsen et al., 2024) consider the model’s own natural language explanation of what it is doing and why. Gradient-based feature attributions have further enabled to scale the localization and analysis of relevant features in LLMs (Ali et al., 2022; Jafari et al., 2024).
@@
Function Vectors and Behavioral Interventions. LLMs’ contextualized representations such as function vectors (Todd et al., 2024) and in-context task vectors (Hendel et al., 2023; Yang et al., 2025) are shown to trigger execution of associated tasks. A broader application of such steering vectors include persona vectors (Chen et al., 2025) and representations of truthfulness (Marks & Tegmark, 2024; Bürger et al., 2024). Probing has identified associations between internal representations and features of interest (Conneau et al., 2018; Hewitt & Manning, 2019). This has also been extended to the analysis of reasoning strategies, e.g., by identifying probes involved in solving grade-school math problems (Ye et al., 2025). These analyses provide some first methods to understand how complex representational geometries and manifolds drive predictions in modern models. Related work from neuroscience shows that task-related variables are encoded in neural geometries in a format that supports generalization to novel situations (Bernardi et al., 2020). Monkey neuroscience research has linked compositional generalization in neural representational geometry with behavioral performance (Fascianelli et al., 2024). These findings suggest that function vector methods can be used to study how compositional generalization may guide algorithmic steering of LLMs, providing a conceptual starting point for our work.
Figure 1: Algorithmic Tracing & Steering: Primitive Extraction & Evaluation. We trace algorithmic primitives by a clustering latent representations, b identifying corresponding reasoning traces, c meta-clustering primitives, identifying sequential transition trends among clusters, and comparing cluster similarity across models and tasks. Once we identify primitives, d we extract associated primitive vectors from top heads, and e use causal patching to validate, explore the compositional geometry and cross-task transfer of primitives.
@@
Algorithmic Evaluation and Steering of Language Models. Multi-step planning and graph navigation are standard benchmarks for structured reasoning (Fatemi et al., 2023), yet most LLMs perform poorly and fail to generalize with growing graph complexity (Momennejad et al., 2023). Although efficient algorithms exist, recent evaluations suggest that LLMs rely on policy-dependent heuristics rather than explicit search strategies (Eberle et al., 2025). Complementary works have aimed to map a high-level causal model to an internal realization by an LLMs (Geiger et al., 2024), and prompt models to generate sets of abstract hypotheses about the tasks to improve inductive reasoning performance (Wang et al., 2024). Recent analysis of chain-of-thought outputs has uncovered sentence-level relationships, offering the localization of salient reasoning steps (Bogdan et al., 2025). Existing work either targets high-level behaviors such as politeness, creativity, or honesty using relatively coarse vector representations to steer behavior (Chen et al., 2025), or focuses on very simple input-output mapping (Todd et al., 2024). Our work focuses on algorithmic steering, targeting computational primitives and their compositions, aiming at a deep understanding of the model’s internal algorithmic phenotypes and applications in future algorithmic finetuning.
@@ 3 Experimental Setup: Models and Tasks
Models. We evaluated decoder-only transformer models on multi-step reasoning tasks. Our analyses primarily focused on Phi-4-Base (Abdin et al., 2024) and the reasoning-specialized variant Phi-4-Reasoning (Abdin et al., 2025), with additional validation from Llama-3-8B (Dubey et al., 2024).
Tasks and Data Collection. Four multi-step reasoning setups were tested. We primarily focused on the Traveling Salesperson Problem (TSP), an NP-hard benchmark. Notably, Phi-4-Reasoning exhibited improved performance on TSP despite not being finetuned for this task (Abdin et al., 2025). We also investigated the 3-SAT and AIME Mathematical Olympiad benchmarks for the same reasons. Finally, we investigated Graph Navigation tasks used in previous LLM reasoning research (Momennejad et al., 2023; Eberle et al., 2025) (see Appendix B for task prompts). Our setups demand planning, search, and verification. TSP and Graph navigation require graph optimization via planning, path generation, search, comparison, and verification. 3SAT presents another NP-hard problem that tests flexible algorithmic problem solving. Finally, AIME requires complex mathematical reasoning and multi-step computation. Together, these setups enable us to examine algorithmic primitives that are task-specific and those that generalize across tasks. We collected reasoning traces and associated residual stream vectors across 100+ examples per task, with systematic variation in problem complexity.
@@ 4.1 Definitions and Notation
Definition: Algorithmic Primitive. We define algorithmic primitive as a minimal computational operation observed in a reasoning process (Eberle et al., 2025) (e.g. TSP), such as retrieving the nearest neighbor, computing a distance, generating a new candidate path, or verifying a solution. Primitives can be identified both in explicit reasoning traces (e.g. reasoning steps the model produces in the output) and in internal activations (e.g. clusters of token representations, or attention patterns).
@@
Definition: Primitive Vector. We define primitive vectors (following the function and steering vector literature (Todd et al., 2024)) as vectors that can be injected into the residual stream to reliably induce a particular primitive. For a given primitive $p$ and transformer layer $\ell$, let the residual stream activation at token $t$ be $h_{\ell}(x,t)\in\mathbb{R}^{d_{\ell}}$. A primitive vector $v^{(p)}_{\ell}\in\mathbb{R}^{d_{\ell}}$ is a direction in activation space that increases the expression of the primitive function $p$. Here we extract primitive vectors using the function vector approach (Todd et al., 2024) (see Section 4.2 for details).
Definition: Primitive Induction for Algorithmic Steering. Given a primitive vector injection strength $\alpha\in\mathbb{R}$, we can intervene on a representation by adding (or patching) a primitive vector:
$\tilde{h}_{\ell}(x,t)=h_{\ell}(x,t)+\alpha\,v^{(p)}_{\ell},$
which increases the probability of expressing $p$ when $\alpha>0$, and decreases it when $\alpha<0$. We refer to this procedure as algorithmic steering or induction.
Algebraic Operations on Primitive Vectors. Assuming compositional geometry, primitive vectors can be combined through simple algebraic operations in activation space. For primitives $p$ and $q$ at layer $\ell$, additive and subtractive composition within a layer (+ and –) can be defined as:
$v^{(p\oplus q)}_{\ell}\approx w_{p}\,v^{(p)}_{\ell}\pm w_{q}\,v^{(q)}_{\ell}.$
Scalar modulation (with varying strength) can be defined as:
$v^{(\alpha p)}_{\ell}=\alpha\,v^{(p)}_{\ell}.$
@@
If this intervention increases the activation
$a^{(p)}_{\ell,T_{2}}(x_{T_{2}},t)=\langle v^{(p)}_{\ell,T_{1}},\,h_{\ell}(x_{T_{2}},t)\rangle$
and increases the associated behavioral hallmark in task $T_{2}$, we denote successful transfer as:
@@
4.2 Algorithmic Tracing & Steering: Primitive Extraction & Evaluation
Step 1. Primitive Identification: Geometric Clustering of Latent Representations. We hypothesize that distinct algorithmic primitives are reflected in higher dissimilarities in the model’s internal representations, and therefore apply k-means clustering (Lloyd, 1982) to the model’s internal representations while processing complex self-generated reasoning traces. For all clustering analyses, we extracted representations from layer 17 of Phi-4-base and used $k=50$ clusters. We fit separate models to TSP, 3-SAT, AIME, and GraphNav and also fit a joint model to TSP+AIME.
Step 2. Primitive Identification: Mapping Clusters to Associated Reasoning Traces. We analyze the primitives associated with the different clusters revealed in step 1 by analyzing the tokens associated with a particular cluster and the context surrounding them. This lets us categorize algorithmic strategies, verification behaviors, and metacognitive patterns and provides direct insight into the reasoning processes models employ.
@@ 4.2 Algorithmic Tracing & Steering: Primitive Extraction & Evaluation
Step 4. Primitive Validation: Primitive Vector Extraction. Finally, we extract primitive vectors associated with particular clusters by adapting the methodology put forward in Todd et al. (2024). We compute the average attention head activations over all tokens that are a part of a particular cluster (using 100 responses) and then average the outputs of the top attention heads that reliably carry out a function ($k=35$; see appendix). This yields a candidate primitive vector for each extracted cluster. We validate candidate primitives by defining behavioral hallmarks associated with their proposed computational role and testing whether injecting the extracted primitive vectors results in an increase in those behavioral hallmarks. Beyond primitive validation, we also examine their compositional generalization by testing their arithmetic composition and cross-task transfer, injecting primitive vectors extracted from AIME into the model while it performs TSP.
Algorithmic Fingerprinting. For a given set of clusters $1,\dotsc,k$ and a response $i$, we extract the relative frequency of tokens assigned to each cluster, $f_{i}\in\mathbb{R}^{k}$, $\sum_{i}f_{i}=1$. We consider $f_{i}$ as a simple “algorithmic fingerprint” of a particular reasoning trace, highlighting differences between the primitives involved in different responses. In particular, we analyze the algorithmic dissimilarity between two responses by computing the symmetric $\chi$-squared distance between their frequencies,
@@
We also consider sets of responses $(f^{(j)})_{j=1}^{n}$, $(g^{(j)})_{j=1}^{m}$. We analyze differences in the involved algorithmic primitives by computing the average frequencies $f=\tfrac{1}{n}\sum_{j=1}^{n}f^{(j)}$, $g=\tfrac{1}{n}\sum_{j=1}^{n}g^{(j)}$ and then computing the signed $\chi$-squared distance
$\chi_{s}(f_{i},g_{i}):=\text{sgn}(f_{i}-g_{i})\circ\frac{(f_{i}-g_{i})^{2}}{f_{i}+g_{i}}\in\mathbb{R}^{k},$ (2)
for each cluster $i$. A large positive difference indicates that the corresponding primitive is much more prominent in the responses $(f^{(j)})_{j=1}^{m}$, a large negative difference indicates that the corresponding primitive is much more prominent in the responses $(g^{(j)})_{j=1}^{m}$. In particular, we compare Phi-4 responses to Phi-4-Reasoning responses and TSP responses to AIME responses.
Figure 3: Clusters are organized in a hierarchical manner and meta-clusters are repeatedly traversed throughout the response. a Cluster-cluster transition matrix on TSP, organized by meta-cluster (MC). Most transitions occur within each meta-cluster. b Average t-SNE trajectories across different highlighted primitives. c Transition between different clusters for an example output. Different meta-clusters are highlighted by colors. The latter half of the response undergoes a cyclical transition. d Most common transitions between different meta-clusters reveals frequently occurring cycles. e-g Cluster-to-cluster transition matrices structured in terms of the inferred meta-clusters on e 3-SAT, f AIME, and g GraphNav.

### v2

@@ Abstract
How do latent and inference time computations enable large language models (LLMs) to solve multi-step reasoning? We introduce a framework for tracing and steering algorithmic primitives that underlie model reasoning. Our approach links reasoning traces to internal activations and evaluates algorithmic primitives by injecting them into residual streams and measuring their effect on reasoning steps and task performance. We consider four benchmarks: Traveling Salesperson Problem (TSP), 3SAT, AIME, and graph navigation. We operationalize primitives by clustering activations and annotating their matched reasoning traces using an automated LLM pipeline. We then apply function vector methods to derive primitive vectors as reusable compositional building blocks of reasoning. Primitive vectors can be combined through addition, subtraction, and scalar operations, revealing a geometric logic in activation space. Cross-task and cross-model evaluations (Phi-4, Phi-4-Reasoning, Llama-3-8B) show both shared and task-specific primitives. Notably, comparing Phi-4 with its reasoning-finetuned variant highlights compositional generalization after finetuning: Phi-4-Reasoning exhibits more systematic use of verification and path-generation primitives. Injecting the associated primitive vectors in Phi-4 induces behavioral hallmarks associated with Phi-4-Reasoning. Together, these findings demonstrate that reasoning in LLMs may be supported by a compositional geometry of algorithmic primitives, that primitives transfer cross-task and cross-model, and that reasoning finetuning strengthens algorithmic generalization across domains.
@@ 1 Introduction
This work aims to understand how fundamental algorithmic primitives are generalized and composed to enable complex reasoning capabilities in LLMs. The work addresses three fundamental questions: (1) What are the basic algorithmic primitives that language models use in specific tasks, and across reasoning domains? (2) Do these primitives compose geometrically in neural activation space? (3) How do reasoning-enhanced models differ from base models in their primitive usage and composition? We introduce a framework for a multi-level algorithmic understanding. We first extract algorithmic primitives by clustering internal representations and leveraging a novel LLM-based automated pipeline for annotating reasoning traces. We then apply function vector methods to extract primitive vectors from the models’ internal representations. We induce and steer algorithmic primitives by injecting primitive vectors across the layers. To identify both task-specific and universal algorithmic primitives, we apply the approach across domains: Traveling Salesperson Problem, 3-SAT, AIME, and graph navigation. Finally, to investigate compositional generalization as a result of finetuning, we compare algorithmic performance of base models and their reasoning-finetuned counterparts, like Phi-4 and Phi4-reasoning (Abdin et al., 2025). This multi-level framework allows us to understand the geometry and compositionality of algorithmic primitives in LLM reasoning.
@@
Figure 1: Algorithmic Tracing & Steering: Primitive Extraction & Evaluation. We trace algorithmic primitives by a clustering latent representations, b identifying corresponding reasoning traces, c annotating them using an automated LLM pipeline, d meta-clustering primitives, identifying sequential transition trends among clusters, and comparing cluster similarity across models and tasks. Once we identify primitives, e we extract associated primitive vectors from top heads, and f use causal patching to validate, explore the compositional geometry and cross-task transfer of primitives.
@@ 2 Related Work
Interpretability for Transformers. With increasing scale and complexity of LLMs, the challenge of understanding their predictions has become an important research direction in explainable AI and interpretability research. This has specifically targeted the analysis of internal model structure and feature relationships (Geiger et al., 2022; Eberle et al., 2022; Schnake et al., 2022), as well as representations and manifolds (Kornblith et al., 2019) including concepts (Chormai et al., 2024). Mechanistic interpretability (Sharkey et al., 2025) has focused on the extraction of circuits (Olah et al., 2020; Wang et al., 2023), feature descriptions (Hernandez et al., 2022), and causally effective representations such as function vectors (Todd et al., 2024). Combined analysis of individual neurons (Gurnee and Tegmark, 2024; Templeton et al., 2024), attention scores (Voita et al., 2019; Clark et al., 2019), and task-specific attention heads (Vig and Belinkov, 2019; McDougall et al., 2024) have further deepened our understanding of internal model processing. In parallel, free-text and chain-of-thought explanations (Turpin et al., 2023; Camburu et al., 2018; Huang et al., 2023; Madsen et al., 2024) consider the model’s own natural language explanation of what it is doing and why. Gradient-based feature attributions have further enabled scaling the localization and analysis of relevant features in LLMs (Ali et al., 2022; Jafari et al., 2024).
@@
Function Vectors and Behavioral Interventions. LLMs’ contextualized representations such as function vectors (Todd et al., 2024) and in-context task vectors (Hendel et al., 2023; Yang et al., 2025) are shown to trigger execution of associated tasks. A broader application of such steering vectors include persona vectors (Chen et al., 2025) and representations of truthfulness (Marks and Tegmark, 2024; bürger2024truthuniversalrobustdetection). Probing has identified associations between internal representations and features of interest (Conneau et al., 2018; Hewitt and Manning, 2019). This has also been extended to the analysis of reasoning strategies, e.g., by identifying probes involved in solving grade-school math problems (Ye et al., 2025). These analyses provide some first methods to understand how complex representational geometries and manifolds drive predictions in modern models. Related work from neuroscience shows that task-related variables are encoded in neural geometries in a format that supports generalization to novel situations (Bernardi et al., 2020). Monkey neuroscience research has linked compositional generalization in neural representational geometry with behavioral performance (Fascianelli et al., 2024). These findings suggest that function vector methods can be used to study how compositional generalization may guide algorithmic steering of LLMs, providing a conceptual starting point for our work.
@@
Algorithmic Evaluation and Steering of Language Models. Multi-step planning and graph navigation are standard benchmarks for structured reasoning (Fatemi et al., 2023), yet most LLMs perform poorly and fail to generalize with growing graph complexity (Momennejad et al., 2023). Although efficient algorithms exist, recent evaluations suggest that LLMs rely on policy-dependent heuristics rather than explicit search strategies (Eberle et al., 2025). Complementary works have aimed to map a high-level causal model to an internal realization by an LLM (Geiger et al., 2024), and prompt models to generate sets of abstract hypotheses about the tasks to improve inductive reasoning performance (Wang et al., 2024). Recent analysis of chain-of-thought outputs has uncovered sentence-level relationships, offering the localization of salient reasoning steps (Bogdan et al., 2025). Existing work either targets high-level behaviors such as politeness, creativity, or honesty using relatively coarse vector representations to steer behavior (Chen et al., 2025), or focuses on very simple input-output mapping (Todd et al., 2024). Our work focuses on algorithmic steering, targeting computational primitives and their compositions, aiming at a deep understanding of the model’s internal algorithmic phenotypes and applications in future algorithmic finetuning.
@@ 3 Experimental Setup: Models and Tasks
Models. We evaluated decoder-only Transformers on multi-step reasoning tasks. Our analyses focused on Phi-4-Base (Abdin et al., 2024) and the reasoning-specialized variant Phi-4-Reasoning (Abdin et al., 2025), with additional validation from Llama-3-8B (Dubey et al., 2024).
Tasks and Data Collection. Four multi-step reasoning setups were tested. We primarily focused on the Traveling Salesperson Problem (TSP), an NP-hard benchmark. Notably, Phi-4-Reasoning exhibited improved performance on TSP despite not being finetuned for this task (Abdin et al., 2025). We also investigated the 3-SAT and AIME Mathematical Olympiad benchmarks for the same reasons. Finally, we investigated Graph Navigation tasks used in previous LLM reasoning research (Momennejad et al., 2023; Eberle et al., 2025) (see App. B for task prompts). Our setups demand planning, search, and verification. TSP and Graph navigation require graph optimization via planning, path generation, search, comparison, and verification. 3-SAT presents another NP-hard problem that tests flexible algorithmic problem solving. Finally, AIME requires complex mathematical reasoning and multi-step computation. Together, these setups enable us to examine algorithmic primitives that are task-specific and those that generalize across tasks. We collected reasoning traces and associated residual stream vectors across 100+ examples per task, with systematic variation in problem complexity.
@@ 4 Methodology
Table 1: Summary of Key Terms and Definitions
Term Definition Example
Algorithmic Primitive A minimal computational operation observed in a reasoning process Retrieving the nearest neighbor in TSP
Algorithmic Tracing The process of identifying all relevant primitives for implementing a particular reasoning process Solving TSP requires distance calculation, nearest-neighbor selection, path comparison, …
Primitive Vector A direction in activation space that can be injected into the residual stream to reliably induce a particular primitive A vector that causes the model to use more nearest-neighbor heuristics when injected
Primitive Induction Modifying model behavior by injecting a primitive vector Injecting a nearest-neighbor primitive vector
@@ 4.1 Definitions and Notation
Definition: Algorithmic Primitive. We define algorithmic primitive as a minimal computational operation observed in a reasoning process (Eberle et al., 2025) (e.g. TSP), such as retrieving the nearest neighbor or computing a distance. Primitives can be identified both in explicit reasoning traces (e.g. reasoning steps the model produces in the output) and in internal activations (e.g. clusters of token representations, or attention patterns).
@@
Definition: Primitive Vector. We define primitive vectors (following the function and steering vector literature (Todd et al., 2024)) as vectors that can be injected into the residual stream to reliably induce a particular primitive. For a given model input $x$, which here refers to the input prompt and the model response, primitive $p$ and transformer layer $\ell$, let the residual stream activation at token $t$ be $h_{\ell}(x,t)\in\mathbb{R}^{d_{\ell}}$. A primitive vector $v^{(p)}_{\ell}\in\mathbb{R}^{d_{\ell}}$ is a direction in activation space that increases the expression of the primitive function $p$. Here we extract primitive vectors using the function vector approach (Todd et al., 2024) (see Section 4.2).
Definition: Primitive Induction for Algorithmic Steering and Causal Evaluation. Given a primitive vector injection strength $\alpha\in\mathbb{R}$, we can intervene on a representation by adding (or patching) a primitive vector: $\tilde{h}_{\ell}(x,t)=h_{\ell}(x,t)+\alpha\,v^{(p)}_{\ell}$, which increases the probability of expressing $p$ when $\alpha>0$, and decreases it when $\alpha<0$. We refer to this procedure as algorithmic steering or induction.
Algebraic Operations on Primitive Vectors. Assuming compositional geometry, primitive vectors can be combined through simple algebraic operations in activation space. For primitives $p$ and $q$ at layer $\ell$, additive and subtractive composition within a layer (+ and –) can be defined as $v^{(p\oplus q)}_{\ell}\approx w_{p}\,v^{(p)}_{\ell}\pm w_{q}\,v^{(q)}_{\ell}$. Scalar modulation (with varying strength) is defined as $v^{(\alpha p)}_{\ell}=\alpha\,v^{(p)}_{\ell}$.
@@
If this intervention increases the activation $a^{(p)}_{\ell,T_{2}}(x_{T_{2}},t)=\langle v^{(p)}_{\ell,T_{1}},\,h_{\ell}(x_{T_{2}},t)\rangle$ and increases the associated behavioral hallmark in task $T_{2}$, we denote successful transfer as:
@@
4.2 Algorithmic Primitive Extraction & Evaluation
Step 1. Primitive Identification: Geometric Clustering of Latent Representations. We hypothesize that distinct algorithmic primitives are reflected in higher dissimilarities in the model’s internal representations, and therefore apply k-means clustering (Lloyd, 1982) to these representations while processing complex self-generated reasoning traces. We chose the k-means algorithm as it provided suitable results on discovering concepts in latent LLM representations in recent investigations (Hawasly et al., 2024; Petukhova et al., 2025; Kopf et al., 2025). For all clustering analyses, we extracted representations from layer 17 of Phi-4-base and used $k=50$ clusters (we justify these choices in App. G). We fit separate models to TSP, 3-SAT, AIME, and GraphNav and also fit a joint model to TSP+AIME.
Step 2. Primitive Identification: Mapping Clusters to Associated Reasoning Traces. We analyze the primitives associated with the different clusters revealed in step 1 by analyzing the tokens associated with a particular cluster and the context surrounding them. This lets us categorize algorithmic strategies, verification behaviors, and metacognitive patterns and provides direct insight into the models’ reasoning processes. To make this process more automatable, we additionally develop an LLM-driven pipeline for automated labeling of the reasoning traces associated with different clusters (App. C). The labels map algorithmic primitives in the reasoning trace to latent clusters. We further define a “cluster expressivity” metric (defined as number of unique tokens expressed by a cluster, divided by the radius of that cluster in latent space, App. A.7), reasoning that clusters with higher expressivity will be more likely to be accurate primitives with clear algorithmic meaning. The LLM-driven pipeline and the cluster expressivity together support directly applying our method to novel models and tasks.
@@ 4.2 Algorithmic Primitive Extraction & Evaluation
Step 4. Primitive Validation: Primitive Vector Extraction. Finally, we extract primitive vectors associated with particular clusters by adapting the methodology put forward in Todd et al. (2024). Specifically, we extract attention head activations for multiple responses and, for each cluster $c$, compute the average activations across all tokens belonging to that cluster. We then average the activations over the top $k=35$ attention heads that reliably carry out in-context learning functions (following the definition in Todd et al. (2024), see App. B.6 for further details).
This yields a candidate primitive vector for each extracted cluster. We validate candidate primitives by defining behavioral hallmarks associated with their proposed computational role and testing whether injecting the extracted primitive vectors results in an increase in those behavioral hallmarks. Beyond validation, we also examine their compositional generalization by testing their arithmetic composition and cross-task transfer, injecting primitive vectors extracted from AIME into the model while it performs TSP.
Figure 2: Clusters are organized in a hierarchical manner and meta-clusters are repeatedly traversed throughout the response. a Cluster-cluster transition matrix on TSP, organized by meta-cluster (MC) (indicated by the white lines). Most transitions occur within each meta-cluster. b Average t-SNE trajectories across different highlighted primitives. c Transition between different clusters for an example output. Different meta-clusters are highlighted by colors. The latter half of the response undergoes a cyclical transition. d Most common transitions between different meta-clusters reveals frequently occurring cycles. e-g Cluster-to-cluster transition matrices structured in terms of the inferred meta-clusters on e 3-SAT, f AIME, and g GraphNav.
Algorithmic Fingerprinting. For a set of $k$ clusters and a response $j$, we extract the relative frequency of tokens assigned to each cluster, $f^{(j)}\in\mathbb{R}^{k}$. $f^{(j)}$ provides an “algorithmic fingerprint”, highlighting differences between the primitives involved in different responses. We measure the algorithmic dissimilarity between responses by the symmetric $\chi$-squared distance between their frequencies,
@@
For sets of responses $(f^{(j)})_{j=1}^{n}$, $(g^{(j)})_{j=1}^{m}$, we measure differences in their algorithmic primitives by computing the average frequencies $f=\tfrac{1}{n}\sum_{j}f^{(j)}$, $g=\tfrac{1}{m}\sum_{j}g^{(j)}$ and then computing the signed $\chi$-squared distance
$\chi_{s}(f_{c},g_{c}):=\text{sgn}(f_{c}-g_{c})\frac{(f_{c}-g_{c})^{2}}{f_{c}+g_{c}},$ (2)
for each cluster $c$. A large positive difference indicates that the corresponding primitive is more prominent in the responses $(f^{(j)})_{j=1}^{m}$, a large negative difference indicates that it is more prominent in the responses $(g^{(j)})_{j=1}^{m}$. In particular, we compare Phi-4 responses to Phi-4-Reasoning responses and TSP responses to AIME responses.